In 1908, Bateson and Punnett (Report to the Evolution Committee of the Royal Society, Report IV), provided data for the F2 progeny from a hybrid experiment in sweet pea (Lathyrus odoratus, which is related to the garden pea, Pisum sativum, which Mendel studied).

BIOL.1001 General Biology 1

Assignment 2

The two traits they studied were flower colour and pollen grain shape. They crossed a true-breeding purple-flowered plant that had long pollen grains with a true-breeding red-flowered plant that had round pollen grains, and tabulated the following results for the F2 progeny:

                                           Purple flowers, long pollen grains        296

                                            Purple flowers, round pollen grains     19

                                            Red flowers, long pollen grains           27

                                            Red flowers, round pollen grains         85

  1. Which traits are dominant and which are recessive? (4 point)
  2. Did the experimental results conform to Mendel’s principle of segregation and independent assortment? Explain why.
  3. Propose an explanation for your finding in 6b.
Answer #1

A)

The dominant traits are purple flowers and long pollen grains.

The recessive traits are red flower and round pollen grains.

Because whn the two parents are crosses together, then there is found long pollen and and purple flowers in F1 progeny These long pollen and purple flowers are dominant traits.

B) No, the experimental results did not confirm to Mendel principle of segregation and independent assortment because in F2 progeny there is is not found equal number of of each individuals. This suggests that there is not found independent assortment of genes and alleles are not segregated properly.

C) There is found linkage behaviour between flower color encoding gene and pollen shape coding genes.

Please look in the image for Cross:

There is found the genetic distance between two gens of:

(Recombinant/total) ×100=(46/427) ×100= 10.77 cM.

That's why there is found recombination frequency between two gene of 10.77% which leads to produce 10.77 % recombinat gametes.